OpenStudy (kenneady):

can SOMEONE PLEASE HELP wht is 2/3+y-4/5=1/3

OpenStudy (wckeller):

I'm assuming we are solving for y. First start out by getting rid of all the fractions. This will make solving this much easier.

OpenStudy (kenneady):

ok then what do i do??

OpenStudy (wckeller):

Multiply everything by 3. Then once you have done that multiply by 5. To get rid of 4/5

OpenStudy (wckeller):

You should end up with: \[2+y-\frac{ 12}{ 5 }=1\] Then multiply by 5: \[10+5y-12=1\]

OpenStudy (kenneady):

oh sorry i was working it out on paper

OpenStudy (kenneady):

and ok thank you!!

OpenStudy (wckeller):

Okay hold on, I believe i have made a mistake

OpenStudy (wckeller):

\[\frac{ 2 }{ 3 } + y-\frac{ 4 }{ 5 }=\frac{ 1 }{ 3 }\] \[y=\frac{ 4 }{ 5 }+\frac{ 1 }{ 3 }-\frac{ 2 }{ 3 }\] Then find the common denominator (15), and multiply the denominator and numerators by 3 (if the denominator is 5) and 5(if the denominator is 3) \[\frac{ 12 }{ 15 }+\frac{ 5 }{ 15 }-\frac{ 10 }{ 15 }\] Final answer is \[y = \frac{ 7 }{ 15 }\]